OpenStudy (anonymous):

-9x+7=3x+19 solve for x?

OpenStudy (anonymous):

x=-1

jimthompson5910 (jim_thompson5910):

-9x+7=3x+19 -9x=3x+19-7 -9x-3x=19-7 -12x=19-7 -12x=12 x=12/(-12) x = -1

OpenStudy (anonymous):

Thank You : ) -TeamShort : )